OpenStudy (anonymous):

Why for sin x + cos x = 0 we cannot convert the sinx to cos x tan x and factorise it out like cos x (tan x + 1) = 0. But for 3 tan x = 2 sin x we can?

ganeshie8 (ganeshie8):

In the second problem, you get solutions : 3sin/cos = 2sin 3sin = 2sin cos sin(3-2cos) = 0 sin = 0, or cos = 3/2 Notice that cos is NOT 0, when u get the other solution.

ganeshie8 (ganeshie8):

For the first problem, you get solutions : sin x + cos x = 0 cos (tan + 1) = 0. cos = 0, or tan = -1 Notice that cos = 0 is a part of solution, so you literally dividing ur given equation by 0, which is illegal. so you cannot divide cos in first problem.

OpenStudy (anonymous):

But for 3 tan x = 2 sin x if we move the 2 sin x to the left then 3 tan x - 2 sin x is equal to zero.

OpenStudy (kc_kennylau):

I wouldn't solve the equation this way.

ganeshie8 (ganeshie8):

yes, so ?

ganeshie8 (ganeshie8):

you need to be careful when u divide a function - u must watch out when the stuff you're dividing both sides becomes 0... otherwise u wil get solutions that make no sense

ganeshie8 (ganeshie8):

3 tan x - 2 sin x = 0 you can divide any function, as long as it doesn't become 0 in the domain of solutions for this equation.

ganeshie8 (ganeshie8):

If you want to divide cos, may solve like below : \(\large \sin x + \cos x = 0\) \(\large \sin x \times \dfrac{\cos x}{\cos x} + \cos x = 0\) \(\color{red}{\cos x \ne 0}\) \(\large \cos x \times \tan x + \cos x = 0\) \(\color{red}{\cos x \ne 0}\) \(\large \cos x (\tan x + 1) = 0\) \(\color{red}{\cos x \ne 0}\) \(\large \cos x = 0, ~ \tan x = -1\) \(\color{red}{\cos x \ne 0}\)

ganeshie8 (ganeshie8):

discard cosx = 0 solution, pick only tanx = -1 part

OpenStudy (anonymous):

Okay ! I got it thanksss!!

ganeshie8 (ganeshie8):

np :) good question ! made me think lol.. thanks for asking too :)